ons er the following information about travelers on vacation: 40% check work erna. 30% use cell phone to stay connected to work, 35% bring a prop with both check work erna and use a cell phone to stay connected, and 54.4% neither check work email ror use a cell phone to stay out of every 100 who bring a laptop also check work email, and 70 out of every 100 who use a cell phone to stay connected also bring a laptop them, 1 m i n 9% ed nor bring a laptop. In addition, 84 (a) What is the probability that a randomly selected traveler who checks work email also uses a cell phone to stay connected? (b) What is the probability that someone who brings a laptop on vacation also uses a cell phone to stay connected? (c) If the randomly selected traveler checked work emalil and brought a laptop, what is the probablity that he/she uses a cell phone to stay connected? (Round your answer to four decimal places.) Need Help?ReTala to Tuter
Answer #1

Solution:

Let E denotes the event that a person checks work email, C denotes the event that the person uses a cell phone to stay connected to work, L denotes the event that the person brings a laptop with him.

We are given:

P(E) = 0.4. P(C) = 0.3. P(L) = 0.35

ΡΕλ C) = 0.19

P(E^{'}cap C^{'}cap L^{'} )=0.544

PEIL) = 0.84

P(LI C.) = 0.7

a) We are required to find:

( EE P ) PGE) = P(

  0.19 0.4

0.475

Therefore, the probability that a randomly selected traveler who checks work email also uses a cell phone to stay connected is 0.475

b) We are required to find:

PlcnL) P(L) P(L|C)P(C) P(L) 0.7 × 0.3 0.6 0.35

Therefore, the probability that someone who brings a laptop on vacation also uses a cell phone to stay connected is 0.6

c) We are required to find here:

P(C|Ecap L)

We are given:

P(E^{'}cap C^{'}cap L^{'} )=0.544Rightarrow P(Ecup Ccup L)^{'}=0.544

P(E, U CU L) = 1-P(EU C U L). = 1-0.544 = 0.456

We know that:

small P(Ecup Ccup L)=P(E)+P(C)+P(L)-P(E|L)P(L)-P(L|C)P(C)-P(Ecap C)+P(Ecap Ccap L)

0.456=0.4+0.3+0.35-0.84 imes 0.35-0.7 imes 0.3-0.19+P(Ecap Ccap L)

0.456=0.356+P(Ecap Ccap L)

small P(Ecap Ccap L)=0.10

P(C|Ecap L)=rac{P(Ecap Ccap L)}{P(E|L)P(L)}=rac{0.1}{0.84 imes 0.35}=0.3401

Therefore, If the randomly selected traveler checked work email and brought a laptop, the probability that he/she uses a cell phone to stay connected is 0.3401